OpenStudy (zenmo):

Need help understanding the solution in Trig. Perform the indicated operation using the standard forms. Round answer to three decimal places.

OpenStudy (zenmo):

\[2\sqrt{2}(\cos(11\pi/12)+isin(11\pi/12)\]

OpenStudy (zenmo):

The solution is -2.732+0.732i. How do you get that?